15. Use the information given below to answer the following questions: U.S. $ Equivalent Currency per U.S. $ Canadian dollar 0.7140 1.4006 British pound 1.6180 0.6180

a) If you have U.S. $750, how many Canadian dollars can you get?

b) If you have 525 British pounds, how many U.S. dollars can you get?

c) How many British pounds can you get for a Canadian dollar? (2 points)
Answer #1

a) 1 U.S. $ = 1.4006 Can. $

Hence, Canadian Dollars in U.S.$750 = 750*1.4006

= Can $1,050.45

b) 1 Pound = U.S. $1.6180

U.S. Dollars in 525 Pounds = 525*1.6180 = USD 849.45

c)British Pounds in 1 Canadian Dollar = (British Pounds/USD)*(USD/Canadian Dollar)

= 0.6180*0.7140

= Pound 0.441252/Can $